The Berkel Corporation manufactures Widgets, Gizmos, and Turnbols from a joint process. June production is 10,000 widgets; 13,500 gizmos; and 15,000 turnbols. Respective per unit selling prices at split-off are $95, $60, and $20. Joint costs up to the split-off point are $187,500. What amount of joint costs will be allocated to the Turnbols? (Do not round any intermediary calculations.)

A) $27,306

B) $17,066

C) $73,726

D) $86,468
Answer #1
Product Output Selling price per unit Sales revenue % of sales revenue Allocated joint cost
Widgets 10,000 95 950,000 950,000/2,060,000 = 46.1165049% 187,500 x 46.1165049% = 86,468
Gizmos 13,500 60 810,000 810,000/2,060,000 =39.3203883% 187,500 x 39.3203883% = 73,726
Turnbols 15,000 20 300,000 300,000/2,060,000 = 14.5631068% 187,500 x 14.5631068% =27,306
$2,060,000 $187,500

The amount of joint costs will be allocated to the Turnbols is $27,306.

Correct option is A.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks